OpenStudy (anonymous):

What equation relates the value of a and b if a is 7 more than the product of 5 and b?

OpenStudy (jlastino):

Take note of the keywords "more than" a is 7 "more than" a= 7 + the product of 5 and b a= 7 + 5b or a= 5b + 7

OpenStudy (anonymous):

could it also be a=7b+5

OpenStudy (jlastino):

No because the problem says "the product of 5 and b"

OpenStudy (anonymous):

Thanks